Solve for x. 3x – 2y = 6
I came up with 2/3y+6. Did I get it?

Answer by praseenakos@yahoo.com(507)

QUESTION:

Solve for x. 3x – 2y = 6


ANSWER:


No. Your answer is not correct....


Here you are asked to solve it for 'x'

For that retain the tern containing x in one side and move all other terms to the other side.



For that add 2y both sides of the expression


==> 3x – 2y + 2y = 6 + 2y


==> 3x = 6 + 2y

Now retain ony x in the left side
( here 3 is multiples by x )

Now divide both sides by 3

==> 3x/3 = ( 6 + 2y )/3


==> x = 6/3 + 2y/3


==> x = 2 + 2y/3 which is the answer.


Hope you found the explanation useful.


Regards.

Praseena.